differece between scalar and dot product
with examples
 

Grade:11

1 Answers

aditya kulkarni
6 years ago
the scalar product of two vectors is the product of magnitude of two vectors and the cosine angle between them
eg: two vectors A&B having magnitude 3 &4 respectively and have angle between them60 degree
the scalar product of the aove two vectors is  ABcos60degree=3*4*0.5=6
 
the vector product of two vectorsis a vector with magnitude equal to the product magnitudes of the two vectors and the smaller sine angle between them and direction perpendicular to the plane containing the two vectors
 
 
eg:vectors A&B with magnitude 5&8 respectively and have angle between them 30 degrees
the scalar product of the two vectors is =ABsin30degree=5*8*0.5n^=20n^    where n^ is a unit vector perpendicular to the plane containing the two vectors
 
 
if you agree with my answer please click “Approve” below